### v1

# MoIRA: Modular Instruction Routing Architecture for Multi-Task Robotics

Abstract
Mixture-of-Experts (MoE) approaches have recently gained traction in robotics applications due to their ability to dynamically allocate computational resources and specialize sub-networks for distinct tasks or environmental contexts, enabling more efficient decision-making. Such systems often comprise sparsely activated experts combined under a single monolithic architecture and require a well-configured internal routing mechanism, which does not allow for selective low-level expert and router customization and requires additional training. We propose MoIRA, an architecture-agnostic modular MoE framework designed to coordinate existing experts with an external text-based router. MoIRA incorporates two zero-shot routing options: embedding-based similarity and prompt-driven language model inference. In our experiments, we choose large Vision-Language-Action models, gr00t-N1 and $\pi_{0}$, as the underlying experts, and train low-rank adapters for low-overhead inference. We evaluate MoIRA on various GR1 Humanoid tasks and LIBERO Spatial and Goal benchmarks, where it consistently outperforms generalist models and competes with other MoE pipelines. Additionally, we analyse the robustness of the proposed approach to the variations of the instructions. While relying solely on textual descriptions of tasks and experts, MoIRA demonstrates the practical viability of modular deployment with precise, low-effort routing and provides an alternative, scalable foundation for future multi-expert robotic systems.
Keywords: Vision-Language-Action models; Modular Policies; Robotic Control; Mixture-of-Experts; Multi-Task Learning; Expert Routing.
1 Introduction
Robotic manipulation and navigation tasks have traditionally been tackled using reinforcement learning (RL) and imitation learning (IL) pipelines. These approaches have demonstrated strong performance across various settings but often depend on dense reward signals, curated expert demonstrations, or extensive task-specific tuning [23, 7, 17]. Transformer-based models such as ACT [30] have further advanced robotic policy learning by enabling fine-grained, sequence-aware control.
More recently, foundation models have emerged as an alternative to traditional RL/IL pipelines, offering general-purpose capabilities without the need for task-specific training or reward engineering. Vision-language models (VLMs) such as PaLI-Gemma [25], LLaVA [19], and Qwen-VL [2] exhibit strong image-text grounding and instruction understanding. Although not designed for robotics, they can interpret natural language commands and scene context, making them useful for high-level planning and zero-shot inference.
Building on this trend, vision-language-action (VLA) models combine vision-language encoders with visiomotor control heads to support end-to-end robotic control. Recent examples include RT-2 [5], the RT-X family [8], OpenVLA [15], MiniVLA [3], $\pi_{0}$ by Physical Intelligence [4], and Gr00t-N1 by NVIDIA [20]. These models are typically pretrained on large-scale, diverse datasets (e.g. Open-X Embodiment [8]) and incorporate heterogeneous data sources, including web-scale multimodal content, subtask annotations, and demonstrations from different robot embodiments. Their goal is to generalize across embodiments, task semantics, and modalities with minimal finetuning. However, their generalist nature can lead to reduced precision, inefficient memory usage, and difficulty scaling to large task libraries [10].
Figure 1: Overview of our MoIRA framework: (1) Fine-tuned agents, each a VLA with a dedicated LoRA adapter; (2) Meta‐descriptions, concise summaries $m_{i}$ characterising each agent’s expertise; and (3) Router core, which infers the specialist index $r$ via one of our two proposed strategies: embedding-based similarity matching (via text encoder $E$) between task $t$ and meta-descriptions $m_{i}$, or prompt-driven language model routing. The resulting index $r$ directs the selected specialist for downstream policy inference.
Concurrently, Mixture-of-Experts (MoE) architectures have regained interest as a modular approach to specialization. Originally proposed by Jacobs et al. [14] for adaptive task decomposition, MoEs are now widely explored for improving efficiency in large language models [6] and in robotic learning [12, 24, 29, 28, 13].
In dexterous manipulation, residual MoE structures [13] enable the composition of expert policies without relying on language inputs or transformers, using fixed routing strategies across object-centric subtasks. MoLe-VLA [28] introduces a spatial-temporal aware router (STAR) to dynamically activate sublayers in VLA models based on token relevance and task structure.
In the locomotion domain, MoRE [29] applies sparse expert activation to quadruped control, routing tasks through a shared vision-language interface to balance sample efficiency with generalization. Similarly, GERM [24] demonstrates expert-based generalist control via MoEs for legged agents.
Additional work explores MoE strategies for policy inference efficiency. MoDE [22] leverages sparse expert denoisers in diffusion transformer policies for LIBERO tasks, while Tra-MoE [27] builds on Any-Point Trajectory Modeling [26] to improve multi-domain policy prediction via sparsely-gated transformers.
While these systems show promise, they share key limitations. Generalist models trained on broad data tend to suffer from degraded per-task performance and high runtime memory usage. Expert-based MoEs require orchestrated training and internal routing mechanisms, often tied to monolithic model structures. This introduces a tradeoff between specialization, modularity, and deployment flexibility that has not been fully addressed.
To address these limitations, we adopt a modular architectural perspective wherein each expert can be independently developed, customized, and optimized. This decouples training from deployment, and allows for flexible reuse across tasks, embodiments, and input modalities. We therefore propose the Modular Instruction Routing Architecture, MoIRA (Figure 1), a framework designed to perform zero-shot model routing via natural language task and expert descriptions. MoIRA sidesteps the scalability constraints of monolithic MoEs by leveraging a pool of pretrained specialists, each finetuned on a focused domain. A lightweight meta-controller dynamically selects the most relevant expert, using either embedding-based similarity or prompt-driven reasoning over the textual task specification.
We evaluate MoIRA on two robotic benchmarks: GR1 [20], which covers embodiment variation (full-body, arms-only, and arms-with-waist), and LIBERO [18], which separates tasks into Goal and Spatial semantic categories. For these experiments, we instantiate MoIRA with the GR00t-N1 and $\pi_{0}$ VLA backbones, using LoRA adapters [11] to enable efficient specialist training. The routing module is pretrained and frozen, requiring no additional tuning to map tasks to experts.
Our contributions are as follows:
  1. 1.
    We propose a novel modular routing architecture, MoIRA, that maps tasks to pre-trained experts based on their textual descriptions.
  2. 2.
    We evaluate two routing strategies – cosine similarity with MiniLM [21]) and prompt-based inference with SmolLM2-1.7B [1] – and demonstrate robustness under perturbed inputs.
  3. 3.
    We validate MoIRA on the GR1 and LIBERO benchmarks, showing that it consistently outperforms generalist models and other MoE approaches on target tasks and previously unseen tasks.
  4. 4.
    We provide an empirical analysis of runtime memory usage and LoRA adapter switching latency, demonstrating the feasibility of multi-expert deployment under hardware constraints.
By decoupling task semantics from execution, MoIRA enables scalable, modular control. It offers a practical step toward robotic agents that generalize across tasks while benefiting from specialization without relying on monolithic training pipelines.
2 Related Work
Classical approaches in robotics rely on RL and IL, both of which have shown strong performance in simulation and real-world environments. RL methods perform well under dense, shaped reward functions, using techniques such as off-policy replay or on-policy actor-critic optimization [23, 7]. However, they often fail in sparse reward regimes due to delayed feedback and exploration bottlenecks. IL methods, particularly behavior cloning [17], bypass reward engineering by mimicking expert demonstrations, yet suffer from distributional shift and limited generalization when faced with unseen states or perturbations.
Foundation models for VLA control. The emergence of foundation vision-language models (VLMs) has shifted the paradigm in robotic perception and planning. Models such as PaLI-Gemma [25], LLaVA [19], and Qwen-VL [2] demonstrate impressive multimodal grounding capabilities, allowing them to parse natural language prompts and interpret visual scenes. While not trained for control, these models have inspired a new class of instruction-conditioned control architectures.
To bridge the gap between perception and action, recent work extends VLMs into vision-language-action (VLA) models. These architectures integrate visual-language encoders with control heads, supporting end-to-end policy learning from natural language instructions. Examples include RT-2 [5] and the RT-X family [8], which are trained on large-scale robot datasets and achieve generalization across tasks. OpenVLA [15] and MiniVLA [3] focus on smaller, instruction-following backbones, while $\pi_{0}$ [4] and GR00t-N1 [20] support scalable deployment through modular and embodiment-aware control heads. These models commonly use datasets like Open-X Embodiment (OXE) [8] and Ego4D [9] to align multimodal inputs with action trajectories.
While VLAs are powerful generalists, their monolithic design often results in inefficiencies when scaled to diverse task domains. They may suffer from overfitting to frequent patterns, struggle with rare or fine-grained instructions, and require extensive resources during inference. These limitations motivate the development of modular alternatives that preserve generalization while enabling efficient task-level adaptation.
Figure 2: Example rollouts from tabletop manipulation tasks using the GR1 Humanoid embodiment and the LIBERO Spatial benchmark with a Panda Franka arm.
Modularity and mixture-of-experts in robotic control. Mixture-of-Experts (MoE) frameworks provide a natural structure for modularization. Initially proposed for adaptive function decomposition [14], MoEs have been extensively adopted in large-scale NLP and are now gaining traction in robotics for task specialization and efficiency [6]. Their appeal lies in sparse expert activation, which allows scaling while keeping inference efficient.
In robotic manipulation, residual MoE architectures [13] demonstrate how multiple base policies can be composed without requiring language conditioning or transformers. These models use fixed routing strategies over handcrafted task splits and achieve strong results in structured environments. MoLe-VLA [28] introduces a learned Spatial-Temporal Aware Router (STAR) that dynamically activates sublayers within a VLA backbone, reducing unnecessary computation while preserving output quality.
For locomotion, Zhao et al. propose MoRE [29], which integrates sparse expert activation into a quadruped control policy using a shared VLM-based interface. Similarly, GERM [24] leverages MoE modules to build scalable, generalist locomotion agents that maintain specialist precision. These works highlight how modularity can be applied across different embodiment types, even in low-level control settings.
Recent work has also explored MoE routing for efficient policy inference. MoDE [22] introduces a diffusion-based transformer model with sparsely activated expert denoisers, targeting sample-efficient multitask learning in manipulation. Tra-MoE [27] builds on Any-point Trajectory Modeling [26], employing a sparsely-gated transformer to improve cross-domain generalization in imitation learning. Both designs rely on internal routing mechanisms and integrated training pipelines, which limit post-hoc adaptability and expert reuse.
Figure 3: Serving and inference configurations in MoIRA. (I) All-agents-in-memory: Each specialist, identified by a unique LoRA adapter ID, is kept resident in GPU memory. This setup enables near-instant agent switching but requires significantly more VRAM, scaling with the number of specialists. (II) Dynamic LoRA adapter loading: A single shared backbone dynamically loads the required LoRA adapter based on router input (e.g., ”Full upper body specialist (ID: 2)”). This configuration reduces memory overhead to one model checkpoint and a single adapter, but introduces up to 9s latency when swapping specialists.
Positioning MoIRA. These prior systems demonstrate the benefits of modularity but also expose critical tradeoffs. Most MoE-based control systems require tightly coupled routing logic, joint training, or customized architectures that reduce interpretability and limit plug-and-play integration. In contrast, MoIRA adopts an external routing design, treating expert selection as an independent classification task driven by pretrained language models. Its use of LoRA adapters enables low-overhead specialization, and its modular layout supports architectural flexibility across tasks and embodiments. By decoupling routing from execution, MoIRA enables expert policies to be updated or replaced independently. Unlike internal token-level gating, its episodic instruction-based routing supports fast switching, simplified deployment, and interpretability of decision boundaries. These design choices make MoIRA especially suited for real-world deployments where modularity, memory efficiency, and runtime specialization are essential.
3 Methods
MoIRA is a modular meta-controller that selects the best-suited specialist policy based on a task description. It is architecture-agnostic, meaning any expert architecture can be used as a backbone, including but not limited to transformer-based or model-based agents. In our work, we focus on foundation VLAs as expert backbones. The specialists are routed based on their textual description. Unlike joint or monolithic models, MoIRA functions as an external routing module, enabling flexible expert integration and efficient deployment.
Figure 4: MoIRA MoE routing module. Given a textual task description and specialist meta descriptions, the router assigns a specialist ID via one of two strategies. A prompt-driven language model (SmolLM2) formats all inputs into a single inference prompt with few-shot examples to infer the matching expert (orange). A lightweight embedding-based method (MiniLM) computes cosine similarity between the task and cached meta descriptions to return the closest expert (purple). Both variants support modular, language-based task routing without requiring direct observation input.
Specialists are routed by MoIRA using embedding-based cosine similarity or prompt-driven language models (Figure 4), with the router requiring no extra training.
At inference time, MoIRA performs two steps: routing and expert execution (Algorithm 1). Given a task instruction $t$ and expert descriptions $\{m_{i}\}_{i=1}^{K}$, it selects the expert index $r$ using either cosine similarity over pretrained text embeddings:
$r\leftarrow\arg\max_{i=1\ldots K}\cos\big{(}E(t),E(m_{i})\big{)},$
or prompt-based reasoning via a frozen LM with Chain-of-Thought formatting:
$r\leftarrow LM_{\text{CoT}}\left([e_{1},\ldots,e_{n}]\parallel t,\{m_{i}\}\right).$
The selected expert $\mathcal{E}_{r}$ is then executed either directly from memory or by loading LoRA adapter weights to produce the output trajectory $\tau$ and result $o$.
Algorithm 1 MoIRA Routing and Execution
1:Task instruction $t$, Expert pool with meta-descriptions $\{m_{i}\}_{i=1}^{K}$
2:Task outcome (trajectory, reward/success signal)
3:Routing:
4:if Embeddings-based routing then
5:     $r\leftarrow\arg\max_{i=1..K}\cos\left(E(t),E(m_{i})\right)$
6:else if Prompt-driven LM routing then
7:     $r\leftarrow LM_{\text{CoT}}\left([e_{1},\ldots,e_{n}]\,||\,t,\{m_{i}\}\right)$
8:end if
9:Expert Execution:
10:if All-agents-in-memory then
11:     Use in-memory model $\mathcal{E}_{r}$
12:else
13:     Load LoRA adapter weights for $\mathcal{E}_{r}$
14:end if
15:Execute task $t$ with expert $\mathcal{E}_{r}$ to produce trajectory $\tau$ and outcome $o$ return $(\tau,o)$
3.1 Benchmarks and Datasets
  1. 1.
    GR1: We use the GR1 benchmark containing three embodiments (arms-only, arms & waist, and full upper body) of GR1 humanoid. Each task contains an annotated task instruction retrieved from tasks.jsonl metadata file.
  2. 2.
    LIBERO: LIBERO benchmark tasks are extracted from .bddl files. We use a modified version of LIBERO splits from [16] and select Spatial and Goal task categories of the benchmark as our semantic axis.
Rollout examples from both benchmarks are visualized in Figure 2.
3.2 Specialist Finetuning
Figure 5: The initial stage of MoIRA’s agentic adapter fine-tuning. We finetune Vision-Language-Action backbones (Gr00t-N1, $\pi_{0}$-base) and derive modular specialists via LoRA: embodiment adapters for GR1 manipulation tasks and semantic adapters for LIBERO goal/spatial tasks. Although demonstrated on VLAs, MoIRA’s unified interface can accommodate any agentic backbone, e.g. transformer-based or model-based approaches.
For GR1, a dedicated GR00t-N1-2B specialist is LoRA fine-tuned for 5K steps for each embodiment. We select three representative tasks to capture variation in embodiment types, task dynamics, and complexity – Pouring, CanSort, and PlacematToPlate. In addition, we train a jointly-tuned generalist model across all three tasks and reserve WineToCabinet (arms & waist), CuttingboardToTieredBasket (arms & waist), and Coffee (full upper body) as previously unseen (hold-out) tasks for evaluation.
For LIBERO, we use the $\pi_{0}$-base-3.3B foundation VLA as the backbone. We LoRA-fine-tune three models: a spatial task expert, a goal task expert, and a generalist, jointly-trained on both tasks – each model tuned for 30K steps using default hyperparameters. The fine-tuning stage is described in Figure 5.
3.3 Routing Strategies
We implement MoIRA task routing using two strategies:
  1. 1.
    Embedding Similarity: We embed both task descriptions and expert meta-descriptions using all-MiniLM-L6-v2 (23M parameters) [21]. Cosine similarity determines the routed expert ID.
  2. 2.
    Prompt-Driven Routing: SmolLM2-1.7B-Instruct [1] receives structured prompts [^1] SmolLM2 receives a prompt containing meta-descriptions of each expert and a short chain-of-thought instruction set. Example tasks with matched specialists are provided to guide classification. The model is then asked to select the most appropriate expert (e.g., Output: 0, 1, or 2) based on task semantics. containing the task and candidate expert descriptions. The model selects the most likely expert via language model inference.
Each expert is annotated with two distinct types of meta descriptions to support routing:
  1. 1.
    Simple: short, literal phrases that describe low-level actions or object interactions using task-specific vocabulary (e.g., picks and places the black bowl).
  2. 2.
    Abstract: higher-level, generalized descriptions that emphasize task intent or spatial reasoning without referencing specific objects (e.g., transports items between spatial zones).
This dual-format setup enables compatibility with both routing strategies: the prompt-driven LM benefits from detailed linguistic cues in abstract descriptions, while the embedding-based method relies on surface-level lexical similarity captured in simple ones. It also allows us to evaluate the router’s generalization behavior across levels of semantic abstraction.
3.4 Classification and Execution
We assign each task a corresponding ground-truth category: embodiment type (for GR1) or semantic type (for LIBERO). We formulate routing as a classification problem. Routing accuracy is measured using the macro-averaged F1 score between predicted and ground-truth labels. At inference time:
  • The router selects the appropriate expert based on the input task description.
  • If using dynamic adapter loading, the corresponding LoRA weights are loaded (swap latency: 9.4s).
  • The selected expert then executes the assigned task.
Swap latency can be amortized over multiple episodes to reduce its impact during extended deployments. MoIRA serving and inference configurations are depicted in Figure 3
3.5 Evaluation Protocol
GR1: We report mean squared error (MSE) of the active joints against the expert observations over 50 evaluation trajectories per task. The main joints are left hand and right hand for full upper body and arms-only, and right hand only for arms & waist embodiment, respectively.
LIBERO: We report Success Rate (SR) over 100 rollouts for subtask category (Spatial and Goal).
We conduct all our experiments on a single NVIDIA RTX A6000 GPU (48GB VRAM). For all training runs, we keep the default hyperparameters and report evaluation results on 5 seeds.
4 Results
For GR1, we analyze specialist finetuning performance, cross-task generalization, instruction robustness, generalization to held-out tasks, and the impact of routing accuracy. For LIBERO, we compare specialist vs. generalist performance, contrast MoIRA with prior MoE systems, assess inference-time efficiency, and outline future directions.
4.1 GR1 Results
Table 1: GR1 Specialists vs. Baseline comparison (MSE ± std over 50 Trajectories)
Model Pouring CanSort PlacematToPlate
Baseline 1.02 ± 0.04 0.70 ± 0.10 1.67 ± 0.52
Specialist 0.008 ± 0.004 0.010 ± 0.010 0.310 ± 0.140
Each task is evaluated using a target-embodiment specialist (3 total).
4.1.1 Specialist Performance vs. Pretrained Baseline
We fine-tune each embodiment-specific specialist for 5K steps and compare it with the pretrained GR00t-N1 baseline in Table 1. We observe significant reductions in MSE: Pouring (1.02 $\rightarrow$ 0.008, 123× reduction), CanSort (0.70 $\rightarrow$ 0.010, 70× reduction), and PlacematToPlate (1.67 $\rightarrow$ 0.31, 5× reduction). These results highlight the effectiveness of embodiment-specific specialization.
Table 2: Various GR1 cross-task MSE ± std evaluation scenarios using two routing types and simple-abstract meta-descriptions
Model Pouring CanSort PlacematToPlate
Baseline 1.023 ± 0.036 0.699 ± 0.091 1.673 ± 0.520
Specialist 0.008 ± 0.004 0.010 ± 0.010 0.311 ± 0.139
Jointly-tuned 0.012 ± 0.009 0.065 ± 0.010 1.187 ± 0.074
MiniLM (simple) 0.009 ± 0.005 0.006 ± 0.005 0.336 ± 0.455
MiniLM (abstract) 0.008 ± 0.002 0.008 ± 0.006 0.798 ± 0.614
SmolLM2 (simple) 0.008 ± 0.006 0.008 ± 0.006 0.241 ± 0.143
SmolLM2 (abstract) 0.009 ± 0.004 0.009 ± 0.007 0.287 ± 0.097
*Each task is evaluated using a target-embodiment specialist (3 total).
4.1.2 Cross-Task Generalization
Table 2 presents all GR1 task evaluations across specialists and the jointly-tuned generalist. While specialists retain near-optimal performance on their respective tasks, the generalist demonstrates broader coverage but lower precision. For example, 1.187 MSE on PlacematToPlate versus 0.311 for the matching specialist.
Figure 6: GR1 Pouring and CanSort perturbed task description MSE with MiniLM router.
Figure 7: GR1 Pouring and CanSort perturbed task description MSE with SmolLM router.
4.1.3 Robustness to Instruction Perturbations
To evaluate routing robustness, we semantically perturb [^2] Perturbed instructions are semantically equivalent rephrasings generated manually for each task, varying lexical choices and syntax (e.g., “pick up the pear on the placemat and set it on the plate” vs “transfer the pear from the mat to the plate”). task instructions by rephrasing them with synonymous verbs and alternate phrasing. For example, ”take the bell pepper from the placemat and move it to the plate” may be rewritten as ”grab the bell pepper off the placemat and put it onto the plate”.
Table 3: Router F1 Score and Held-out GR1 Tasks Evaluation (MSE $\pm$ std over 10 episodes per task)
Model WineToCabinet Cuttingboard Coffee F1 (avg)
MiniLM1 1.669 ± 0.047 0.399 ± 0.287 0.201 ± 0.023 1.00
MiniLM2 1.662 ± 0.046 0.294 ± 0.209 0.196 ± 0.018 1.00
SmolLM21 1.669 ± 0.046 0.735 ± 0.677 0.192 ± 0.019 0.87
SmolLM22 1.664 ± 0.047 0.482 ± 0.413 0.197 ± 0.016 0.97
CuttingboardToTieredBasket.
1 abstract meta descriptions; 2 simple meta descriptions
As shown in Table 4 and Figures 6-7, SmolLM2 maintains stable performance across tasks regardless of prompt type. In contrast, MiniLM (abstract) shows sharp degradation (e.g. MSE=0.16 on CanSort), indicating higher sensitivity to linguistic variation.
4.1.4 Generalization to Held-Out Tasks
We evaluate MoIRA on three held-out GR1 tasks. As shown in Table 3, MiniLM (simple) achieves perfect routing (F1 = 1.00) with low MSE across tasks. In contrast, SmolLM2 (abstract) yields lower MSE on CuttingboardToTieredBasket (0.735) but suffers from reduced routing reliability (F1 = 0.87). MoIRA significantly outperforms a jointly-tuned generalist model on two out of three tasks, achieving lower MSE on WineToCabinet (1.55 vs. 1.87) and CuttingboardToTieredBasket (0.41 vs. 1.29), as shown in Table 5. These results support the effectiveness of modular specialization when paired with accurate routing.
Table 4: Router MSE on Original vs. Perturbed Task Descriptions (10 Episodes per Task)
Routing Pouring CanSort Placemat
Orig. Pert. Orig. Pert. Orig. Pert.
MiniLM simple 0.009 0.008 0.006 0.009 0.336 0.423
MiniLM abstract 0.008 0.008 0.008 0.160 0.798 0.412
SmolLM2 simple 0.008 0.005 0.008 0.009 0.241 0.293
SmolLM2 abstract 0.008 0.007 0.009 0.008 0.287 0.317
4.1.5 Routing Accuracy and Control Performance
Accurate routing is essential for effective control. When the routing F1 score is near 1.0, all models yield low MSE. However, control quality deteriorates once routing F1 drops below $\sim$0.9. This trend is most evident in MiniLM (abstract), which frequently misroutes tasks under perturbation. In contrast, SmolLM2, particularly with simple prompts, maintains strong routing fidelity and downstream control performance due to more reliable classification.
Table 5: MoIRA vs. Jointly-Tuned Generalist Performance on Held-out GR1 Tasks (MSE $\pm$ std over 10 episodes per task)
Model Wine Cuttingboard Coffee
MoIRA 1.55 $\pm$ 0.04 0.41 $\pm$ 0.26 0.19 $\pm$ 0.02
Joint Generalist 1.87 $\pm$ 0.05 1.29 $\pm$ 0.06 0.16 $\pm$ 0.01
CuttingboardToTieredBasket.
4.2 LIBERO Results
4.2.1 Specialist vs Generalist Success Rates
Figure 8 and Table 7 report Success Rates (SR) on the LIBERO Spatial and Goal subsets. The baseline $\pi_{0}$ model without task-specific fine-tuning achieves 0% SR, as it lacks the necessary observation normalization statistics required for evaluation on these tasks.
Table 6: Comparison of MoE Flexibility and Success Rates (SR) on the LIBERO Benchmark
Method Spatial Goal Routing
Tra-MoE + mask 73 78 Internal
MoDE 90 97 Internal
MoIRA (ours) 94 93 External
In contrast, specialists reach 94% (Spatial) and 93% (Goal), while the jointly-tuned generalist achieves 95% and 90%, respectively. These results highlight the importance of correct routing under MoE settings for performing on par or better than a generalist model.
4.2.2 Comparison to Prior MoE Systems
Table 6 compares MoIRA to Tra-MoE and MoDE. MoIRA’s external routing achieves competitive SRs while retaining modularity and expert flexibility. In contrast to internal routing (e.g., token-level gating in MoDE), MoIRA’s design enables episodic routing with adaptable backbones.
Figure 8: Success Rate (SR) comparison on LIBERO subtasks using the $\pi_{0}$-base model. We evaluate a fully finetuned generalist (gray), jointly-trained LoRA expert on spatial and goal tasks (green), and individual LoRA specialists for each subtask (blue and yellow). All models are trained for 30k steps with default hyperparameters.
Reuss et al. [22] achieve 90% SR on LIBERO Spatial and 97% LIBERO Goal using a sparse expert diffusion transformer. In contrast, MoIRA achieves 93.5% weighted-average SR across Spatial and Goal subtasks using modular $\pi_{0}$ specialists routed via a small language model. Unlike MoDE’s internal token-level routing, MoIRA supports episodic routing via fast adapters, making it suitable for deployment-limited settings.
Table 7: LIBERO Specialists and Generalist Success Rates over 100 Rollouts
Model LIBERO Spatial, % LIBERO Goal, %
Spatial Specialist 94
Goal Specialist 93
Joint Generalist 95 90
We also compare our approach with Tra-MoE [27], which achieves 73% SR on LIBERO-Spatial and 78% on LIBERO-Goal using a transformer-based policy with internal MoE blocks. MoIRA outperforms Tra-MoE on both suites while offering broader routing flexibility and simpler adapter management. These results underscore MoIRA’s strength in maintaining task specialization while scaling across domains, outperforming both internal MoE and diffusion-based approaches in success rate and deployment modularity.
4.2.3 Inference Efficiency and Adapter Management
MoIRA performs runtime specialist switching by loading lightweight LoRA adapters. When all adapters are preloaded in GPU memory, switching between experts is effectively instantaneous, but GPU memory usage scales linearly with the number of specialists. Alternatively, dynamic loading of adapters from CPU reduces memory demands to a single backbone and one active adapter but introduces a latency of approximately 9 seconds per swap (for VLA backbones).
The dynamic setup is especially suitable for amortized, asynchronous use cases, such as domestic service robots completing batches of spatial tasks before switching to goal-directed behaviors.
4.2.4 Routing Fidelity and System-Level Gains
Consistent with GR1 findings, routing accuracy in LIBERO strongly correlates with downstream task success. SmolLM2 achieves perfect classification performance (F1=1.0) in a zero-shot manner, enabling specialist experts to reliably outperform the jointly-tuned generalist. These results underscore that routing fidelity is the critical point for realizing the full potential of modular expert systems.
5 Discussion
This work presents MoIRA, a modular MoE routing framework for robotic manipulation tasks. It dynamically assigns pretrained VLA specialists based on robot embodiment or task semantics. To our knowledge, it is the first framework to demonstrate benchmark-validated dynamic specialization of VLA models along these axes. MoIRA uses LoRA adapters and natural language routing to enable strong end-to-end control without requiring MoE retraining. Our results show that accurate routing, rather than further finetuning, is key for the performance gains observed with MoIRA. Evaluation across GR1 and LIBERO benchmarks confirms that task-specific experts, when routed correctly.
The system offers several practical benefits. By finetuning specialists on narrow task distributions, MoIRA improves downstream performance while maintaining modularity. Since experts are decoupled, they can be updated independently, enabling incremental development and domain-specific adaptation. Instruction-following models such as SmolLM2 provide reliable zero-shot routing even under perturbed natural language inputs. MoIRA’s modularity enables per-expert optimization, such as distillation or quantization, without modifying the overall architecture. This supports efficient deployment on low-resource agents while preserving specialization and routing flexibility.
Still, MoIRA has several limitations. Dynamic adapter swapping introduces a latency of $\sim$9 seconds, which limits responsiveness in interactive settings. Preloading all adapters into GPU memory mitigates this at the cost of linearly increasing VRAM usage. All evaluations are currently in simulation. Real-world deployment would introduce additional variability through actuation noise, latency, and sensory drift, which may challenge routing stability. Model performance also depends on router accuracy: while SmolLM2 is robust, MiniLM is sensitive to phrasing and ambiguity. The system currently lacks fallback logic; misclassifications cannot be corrected by reverting to a generalist. Additionally, expert meta-descriptions are handcrafted, which creates manual overhead and limits scalability across new task sets.
Another core limitation is MoIRA’s reliance on textual task descriptions for routing. While this design aligns with user-issued commands, it assumes well-formed instructions and ignores contextual signals such as visual observations or environment state. Future work could explore multimodal routing approaches that incorporate perception, history, or task grounding to improve robustness. Real-world validation under physical constraints is a key direction. Incorporating confidence-aware fallback routing could mitigate failures caused by misclassification. Techniques such as adapter caching, speculative routing, and concurrent model loading may reduce latency and improve responsiveness. Another promising avenue is reusing the pretrained language encoder from the VLA backbone as the router, which could eliminate the need for separate models and streamline execution. Finally, while MoIRA relies on manually written meta-descriptions, automating this step using language models could unlock fully autonomous deployment. Generating expert summaries from raw task specs, sensor data, or demonstrations would remove manual bottlenecks and make instruction routing fully end-to-end.
In summary, MoIRA offers a modular, scalable approach to deploying robotic agents that combine embodiment-specific specialization with semantic task understanding. It demonstrates that pretrained experts can be orchestrated through zero-shot lightweight text-description-based routing, enabling generalization without relying on joint training or monolithic models.